### v1

# Mapping Overlaps in Benchmarks through Perplexity in the Wild

Abstract
We develop signatures of capacity familiarity to characterize large language model (LLM) benchmarks and their meaningful overlaps. Benchmark signatures probe the capacity required for benchmark performance. We formally define them as a set of salient tokens drawn from in-the-wild, naturally authored corpora, where LLM token perplexity, reflecting more or less pre-training exposure, becomes highly predictive of LLM benchmark performance. Through a large-scale meta-evaluation, we extract benchmark signatures via stepwise forward selection with linear regressions across 32 LLMs and 88 benchmarks spanning diverse knowledge, coding, logic, instruction following, math, language, reasoning, and cultural/world modeling. Our analysis situates signatures in relation to both the semantic similarity of benchmark questions and the correlation of model performance. While performance overlaps are universally high and semantic overlaps remain confined to a narrow mid-range, benchmark signatures prove highly informative in capturing variation, overlap, and divergence. We observe overlap in knowledge and reasoning subtasks, whereas multilingual and cultural benchmarks exhibit less similarity, even compared to cross-task overlap. Notably, performance-level results are strongly influenced by benchmark-orthogonal factors such as question format, highlighting limitations in LLM generalization, the conflation of performance with ability, and issues inherent in current mainstream benchmark agreement studies. Benchmark signatures, however, remain robust to such effects. Ultimately, we identify cross-functional overlaps across logic, math, language, instruction following, and world modeling, with coding emerging as the least overlapping domain. Together, these findings provide mechanistic insights into benchmark validity and LLM sensitivities, and sketch the underlying landscape of interconnected LLM capabilities.
Figure 1: Left: Signature-based correlations across benchmark functions. The Spearman correlation is on average 0.285 for benchmarks within the same design function and 0.087 for cross-function overlaps. Right: Performance-level correlations grouped by benchmark families (MMLU vs. BBH) and question formats (Multi-Choice Questions vs. True-or-False). Mainstream performance-based benchmark agreements are biased towards these benchmark-orthogonal factors (red areas) rather than actual design functions. $\rho_{s}$ represents the similarity range in the right panel.
1 Introduction
Benchmarks have been central in the growth of large language models (LLMs): they catalyze progress, standardize evaluation, and enable systematic cross-model comparisons, thereby influencing the trajectory of AI research. The community has witnessed an accelerating proliferation of benchmarks across a wide range of LLM abilities, such as reasoning (Tafjord et al., 2020) and agentic capabilities (Zhu et al., 2025), as well as real-world scenarios such as finance (Zhang et al., 2023) and safety (Mou et al., 2024). The dedicated “Datasets and Benchmarks” track in leading venues such as NeurIPS and KDD highlights both the importance and steady growth of this area. Each year witnesses many new benchmark papers. From 252 submissions to the NeurIPS Datasets and Benchmarks Track in 2021 to 1,820 in 2024  [^1] https://papercopilot.com/, the number of benchmark papers has increased more than sevenfold. While these resources often claim to assess distinct capabilities, it is frequently unclear whether they truly do so, or whether they merely capture narrow proxies, prompt-specific heuristics, or even overlapping skills that have already been extensively tested elsewhere, making them less unique and useful than advertised. This raises critical questions: Do we really need such a vast and ever-expanding suite of benchmarks? How much overlap exists across them? Answering this question will also reveal the converse: What areas of capability are sparsely underrepresented by benchmarks and might benefit from more?
In this paper, we undertake a comprehensive meta-evaluation with a particular focus on identifying and analyzing benchmark overlap, which we define as the degree to which two benchmarks evaluate a shared set of model capabilities. To capture overlap in a principled way, we examine it from three complementary perspectives. At the semantic level, we assess whether the questions in two benchmarks substantially overlap in content or intent; if so, their redundancy is intrinsic. At the performance level, the mainstream level in benchmark agreement studies (Perlitz et al., 2024), we test whether models show highly correlated performance across two benchmarks, indicating that they measure related underlying abilities even if under surface semantic differences. Finally, at the benchmark signature level - introduced by us in Section 3 - we move beyond tasks and outcomes to characterize the distributional fingerprint of benchmarks, defined by token-level perplexity patterns on large-scale in-the-wild corpora.
Why do in-the-wild corpora effectively encode benchmark characteristics? The abilities measured by benchmarks - commonsense, factual memory, scientific reasoning, programming skills, and more - do not emerge out of thin air. They stem from the diverse real-world text patterns encountered by the model during pretraining. In-the-wild corpora, consisting of large-scale, naturally authored, multi-domain text and code (news, forums, encyclopedias, textbooks and notes, papers, documentation, blogs, and repositories), are produced for human communication rather than adapted for benchmark design. They are rich in task-bearing structure (question–answer, problem–solution, claim–evidence, instruction–execution), redundancy (the same function expressed in many ways), and breadth. This breadth of distribution - likely unique to in-the-wild data - forms the “soil” from which such capabilities grow, and also the source from which benchmark questions are drawn. Even if a benchmark item never appears verbatim, its “function” recurs pervasively: unit-aware arithmetic in recipes (“double 1½ cups”), commonsense causality in narratives (“the glass shattered after being dropped”), claim $\rightarrow$ measurement $\rightarrow$ inference chains in scientific abstracts, code repair patterns in GitHub issues (“off-by-one in loop; fix bounds”), and even schema–query mappings (“customers with orders in last 30 days”). Focusing only on synthetic or benchmark-adjacent data risks capturing artifacts of test design. In-the-wild data, by contrast, mirrors the true distribution that gives rise to these abilities, making the overlap between pretraining exposure and benchmark competence not accidental but expected.
Perplexity provides a useful lens for quantifying relationships between pretraining, skills, and benchmark performance. Low perplexity on a passage suggests that the model has seen similar linguistic and conceptual patterns during training and is familiar with the content. High perplexity, by contrast, indicates unfamiliarity and underrepresentation in pretraining. Thus, the distribution of perplexity values across large corpora serves as a fingerprint of the model’s training exposure and more or less acquired capacity. Importantly, because different benchmarks stress different capabilities, they map onto different perplexity distributions when probing across the same corpus. In other words, corpora encode benchmark signatures because benchmarks are not foreign entities imposed on the model after training, but rather structured samplings of capabilities that themselves emerge from the distribution of in-the-wild data. Perplexity serves as the bridge between corpus exposure and benchmark performance, making it possible to identify and characterize these signatures without requiring direct evaluation on the benchmark itself  [^2] More discussions about prior works see A.1.. We therefore leverage perplexity as the basis and covariate for salient token selection and signature formation  [^3] More details see Section 3 and Appendix A.2.. The following three levels in this work provide a holistic framework: semantics address task design, performance captures model behavior, and signatures reveal a fingerprint of model capacity. The overlap between benchmarks across each of these levels highlights the interconnected capacity space - an oft-discussed yet difficult-to-formalize concept and so represents a promising tool for evaluating benchmark validity. This rationale is illustrated in Figure 2.
Figure 2: Overview of the rationale of how in-the-wild corpora implicitly encode the benchmark signature, knowledge exposure during pretraining (capacities), as well as benchmark performance.
Definition: Benchmark Signature A benchmark signature is defined as a set of salient tokens $T$, extracted from large-scale in-the-wild corpora, such that the perplexity of a collection of language models $M$ on $T$ is highly predictive of their performance on the benchmark.
To achieve the overall process, we make the following three contributions:
  • We introduce a systematic framework for measuring benchmark relations and especially their overlap across three levels: semantic, performance, and signature derived from model perplexity.
  • We develop a forward selection and regression-based pipeline to extract these signatures by mining and filtering token-level perplexity statistics from in-the-wild pretraining corpora.
  • We uncover unexpected overlaps between widely used benchmarks. While these benchmarks are intended to test a specific ability - such as logic - and their problem sets do align with human intuitions about logic, in practice they often measure instruction-following ability in language models instead. This reveals the potential issue of benchmark design and actual execution, as well as the interconnected space of LLM capabilities.
2 Semantic Overlaps and Performance Overlaps
General Notation.
We denote the collection of $m$ LLMs by $M_{1},\ldots,M_{m}$} and the set of $n$ benchmarks by $B_{1},\ldots,B_{n}$}. For any quantity defined jointly over a model-benchmark pair—such as a performance metric $y$ - we write $y_{i,j}$ to indicate the metric value corresponding to model $M_{i}$ evaluated on benchmark $B_{j}$. Unless otherwise specified, all vectors are column vectors and are set in bold lowercase, e.g. $\mathbf{x}\in\mathbb{R}^{d}$. Matrices are represented with capital letters in bold, e.g. $\mathbf{X}\in\mathbb{R}^{n\times m}$.
Semantic-Level Overlap.
For benchmarks $B_{a},B_{b}$ with question text sets $Q_{a},Q_{b}$, let $n_{\min}=\min\{|Q_{a}|,|Q_{b}|\}$. Let $k$ be the embedding dimension, $f:\text{text}\to\mathbb{R}^{k}$ be a sentence transformer (e.g. MPNet encoder; Song et al. (2020)), and let $s(x,y)$ be the cosine similarity between $x,y\in\mathbb{R}^{k}$. Because benchmarks vary in size (i.e., number of questions), which could bias results, we estimate overlap via size-matched bootstrapping similarity: for $T=1000$ times, we draw $\{q\}_{t}^{(a)}\subset Q_{a}$ and $\{q\}_{t}^{(b)}\subset Q_{b}$ independently with $|\{q\}_{t}^{(a)}|=|\{q\}_{t}^{(b)}|=n_{\min}$, encode each item with $f$. Also, $Concate()$ means concatenating a list of texts into a single string within each set, and computing cosine similarity:
$\widehat{A}_{\mathrm{sem}}(B_{a},B_{b})=\frac{1}{T}\sum_{t=1}^{T}s\bigl(f(concat(\{q\}_{t}^{(a)}),f(concat(\{q\}_{t}^{(b)}))$
Overall, this mitigates sample-size bias and yields a more robust similarity estimate. Full procedural details appear in Appendix A.4.1. The overlap between $B_{a}$ and $B_{b}$ is defined by the $\widehat{A}_{\mathrm{sem}}(B_{a},B_{b})$.
Performance-Level Overlap.
For each benchmark $B_{a}$, let $\mathbf{y}_{:,a}\in\mathbb{R}^{m}$ be the vector of model performances on $B_{a}$ (one entry per model). The performance-level overlap between two benchmarks $B_{a}$ and $B_{b}$ is the Spearman rank correlation between their model-marginalized performance vectors:
$\rho(B_{a},B_{b})\;=\;\operatorname{corr}\!\bigl(\operatorname{rank}(\mathbf{y}_{:,a}),\,\operatorname{rank}(\mathbf{y}_{:,b})\bigr)$
Thus, the overlap between $B_{a}$ and $B_{b}$ under performance-level is defined by the spearman correlation which is $\rho(B_{a},B_{b})$.
3 Mining Benchmark Signatures from In-the-wild data
Algorithm 1 Obtaining signature for benchmark $B_{j}$
0:  Data “in the wild” $\mathcal{D}$, Benchmark $B_{j}$, a list of LLMs $M_{1},...,M_{m}$
0:  Signature $S_{j}$
1:  $y_{:,j}\leftarrow B_{j}$ with $M_{1},...,M_{m}$ $\triangleright$ Generate performance column vector on benchmark $j$.
2:  $\mathcal{T}\leftarrow\mathcal{D}$ $\triangleright$ Processing in-the-wild data to tokens with preceding context.
3:  $\mathbf{P}\leftarrow\mathcal{T}$ With $M_{1},...,M_{m}$ $\triangleright$ Generate the token-level perplexity covariate matrix.
4:  $\mathcal{T}^{\prime}_{j}\leftarrow\textsc{AIC}\!\left(\textsc{ThrushPrefilter}(\mathbf{P}\sim y_{:,j})\right)$ $\triangleright$ Perform Thrush pre-filtering first; then stepwise AIC feature selection on the covariate matrix $\mathbf{P}$ against the performance vector $y_{:,j}$ of benchmark $B_{j}$ to obtain salient tokens.
5:  Retrieve $S_{j}$ from mapping $\mathcal{T^{\prime}}_{j}$ in $\mathbf{P}$
6:  return $S_{j}$
The overall process of mining signatures can be found in Algorithm 1 and its details can be found in Appendix 4. Let $d$ denote the number of “in-the-wild” tokens (in our case, tokens drawn from large-scale pretraining corpora [^4] Progressing from fine to coarse granularity, we have token-, chunk-, and document-level perplexities. We provide more experimental results of why the token-level operation is the best. Details are shown in Appendix A.2.), denoted as $\mathcal{T}=\{t_{1},\dots,t_{d}\}$, where $d$ typically scales to billions. For any benchmark $B_{j}$, our objective in extracting its benchmark signature is to isolate a subset of salient tokens $\mathcal{T}_{j}^{\prime}=\{t^{\prime}_{1},\dots,t^{\prime}_{d^{\prime}}\}\subset\mathcal{T}$ that are maximally informative in explaining variations in LLM performance. We formalize this as a regression problem: let $\mathbf{\hat{y}}_{j}:=(y_{1,j},\dots,y_{m,j})^{\top}\in\mathbb{R}^{m}$ denote the performance of $m$ language models on benchmark $B_{j}$. The covariate matrix $\mathbf{P}\in\mathbb{R}^{m\times d}$ contains token-level perplexities, where entry $\mathbf{P}_{ij}\equiv p_{ij}$ corresponds to the perplexity of token $t_{j}$ under LLM $M_{i}$. The challenge lies in the high-dimensional regime ($d\gg m$), where classical regression approaches are ill-posed. To make progress, we must uncover and exploit latent structural properties of the problem. In particular, we put forward a key assumption and a follow-up question:
  1. 1.
    Sparsity: Most token-level perplexities are uninformative for predicting benchmark performance, with only a small fraction carrying predictive signals. [Assumption 1]
  2. 2.
    Extraction: Assuming sparsity holds, what methods can effectively identify and extract these predictive signals from the overwhelming background of noise? [Question 1]
Together, they motivate our below regression-based framework for mining benchmark signatures, which leverages high-dimensional inference techniques to disentangle signal from noise and recover benchmark-specific fingerprints of token-level perplexity.
3.1 Token-level Filtering with Perplexity Correlations
Figure 3: Distribution of Thrush correlations in pre-selection phases; red vertical lines mark the 1st and 99th percentiles, highlighting that few features are highly correlated with performance.
Answering [Q1] by fitting a full multivariate regression model is computationally intractable given that the number of tokens $(d)$ is orders of magnitude larger than the number of models $(m)$. We therefore adopt a pragmatic and efficient two-stage approach, beginning with a screening step to drastically reduce the feature space. Specifically, for each benchmark, we perform a token-by-token correlation screening. We compute a robust correlation coefficient between each token’s perplexity vector and the benchmark performance vector. This screening is highly efficient, requiring linear time in the number of features, $O(md)$, and allows us to observe the empirical distribution of these coefficients. In a sparse regime, we expect this distribution to be sharply peaked at zero, with small tails representing potentially informative tokens.
A known limitation of this screening approach is its reliance on marginal, univariate correlations. It evaluates each token in isolation, potentially overlooking features that are predictive only in a multivariate context (e.g., suppressor tokens that explain residual variance). However, we argue this approach is theoretically and empirically well-justified in our specific problem setting for the following reasons:
  1. 1.
    Justification from the Ultra-High Dimensional Regime: Our problem, with $d\gg m$, resides in the ultra-high dimensional setting. Theoretical frameworks developed for this regime, such as Sure Independence Screening (SIS; Fan & Lv (2008)), provide formal guarantees for marginal screening. The “sure screening property” ensures that, under sparsity and certain regularity conditions, correlation-based filtering can discard the vast majority of irrelevant features while retaining the true predictive signals with very high probability. We further explain how several key conditions of SIS are plausible in our context in Section A.3.
  2. 2.
    Empirical Precedent in Data Selection: This screening methodology has demonstrated strong empirical success in the related domain of data selection for pre-training. Prior work has successfully used document-level perplexity correlations to filter large corpora, improving downstream model performance(Thrush et al., 2025; Shum et al., 2025). Their success provides compelling evidence for the practical utility of correlation screening as a robust heuristic for identifying informative signals in LLM-related data.
While there exist various methods for robust correlation calculation, there is no single “silver bullet”; the choice is often guided by the specific properties of the data. In particular, we highlight the two robust correlation coefficients introduced in the aforementioned data selection literatures.
Definition 3.1 (Thrush Correlation (Thrush et al., 2025)).
Fixing the $j$-th token $t_{j}\in\mathcal{T}$. Let $\text{rank}_{j}(p)$ denotes the rank of $p$ among $\{p_{1,j},\cdots,p_{m,j}\}$ and $\text{sign}(\cdot)$ be the sign function, we denote
$\gamma_{j}=\sum_{1\leq k<l\leq m}\text{sign}(y_{k,j}-y_{l,j})(\text{rank}_{j}(p_{k,j})-\text{rank}_{j}(p_{l,j}))$ (1)
as the Thrush correlation coefficient. This coefficient is a variant of Kendall’s $\tau$ (Kendall, 1938), measuring the concordance between model performance and perplexity ranks. It counts the number of model pairs where the model with better performance also has a lower perplexity rank (a concordant pair), and subtracts the number of pairs where this is not the case (a discordant pair), making it robust to the absolute magnitude of perplexity values.
Definition 3.2 (Pre-select Correlation (Shum et al., 2025)).
Letting $Z=\frac{m(m-1)}{2}$ to be a normalizing factor, and (1),,(m)(1),...,(m) be the sorted indices by LLM performances (i.e. $y_{(1),j}\leq y_{(2),j}\leq\cdots\leq y_{(m),j}$), the Pre-select correlation coefficient is defined as:
$\eta_{j}=\sum_{1\leq k<l\leq m}\mathbf{1}\{p_{k,j}>p_{l,j}\}/Z$ (2)
The Pre-select coefficient computes the fraction of model pairs that are “misordered” by their token perplexities relative to their benchmark performance. In an ideal scenario where lower perplexity perfectly predicts higher performance, this sum would be zero; a value of 0.5 would indicate a random, uninformative relationship.
Once these robust correlation coefficients are calculated for all $d$ tokens, we employ a simple quantile-based threshold to screen the feature space, retaining approximately the top 1% of tokens with the strongest signal. Figure 3 presents the empirical distributions of the Thrush coefficients for three representative benchmarks. In all cases, the distributions are sharply peaked around a central value (indicating a random relationship), with thin tails representing tokens that are highly correlated with performance. This characteristic shape provides compelling empirical support for our sparsity hypothesis ([Q1]): the vast majority of token perplexities are uninformative, while a small, identifiable subset carries a significant predictive signal.
3.2 Refining signatures with forward selection regression
The correlation screening successfully isolates a candidate set of potentially informative tokens, satisfying our goal of drastically reducing the search space. However, this filtering alone is insufficient to define a robust benchmark signature for two primary reasons. First, the filtered set is likely to contain redundant features; for instance, several top-ranked tokens might represent the same underlying linguistic phenomenon and thus offer overlapping predictive information. Second, a true signature should not only identify important tokens but also capture their conditional importance – their predictive power given the other tokens already in the model.
To address these challenges and distill a final, parsimonious signature, we employ a second-stage multivariate variable selection procedure. Our general framework can accommodate various high-dimensional regression techniques suited for the $d^{\prime}>m$ regime (where $d^{\prime}$ is the number of filtered tokens), including penalized methods like Lasso (Tibshirani, 1996), Ridge (Hoerl & Kennard, 1970), or Elastic Net (Zou & Hastie, 2005). In practice, we opt for a greedy forward selection approach, which we find builds interpretable and effective models. This method iteratively constructs the signature by adding the single token from the candidate pool that yields the greatest improvement to the model’s fit, penalized by its added complexity.
To guide this selection process, we use the Akaike Information Criterion (AIC; Bozdogan (1987)), which provides a principled trade-off between explanatory power and model size, mitigating the risk of overfitting. The process terminates when no additional token can improve the model’s AIC score by a meaningful amount. The complete two-stage process – combining the initial correlation screening to create a candidate set with the subsequent forward selection to derive the final signature – is formalized in Algorithm 4.
3.3 Signature-Level Overlap
Consider two benchmark signature vectors, $S_{1}$ and $S_{2}$, each including several pieces of context (30 pieces separated by space) + the salient token. We use 32 models to process these signatures, reading their respective pre-contexts, producing the last token-level perplexities and calculating overlaps. If the models are confused to a similar degree by both signatures, that is a strong indicator that the two benchmarks align. Since some “weak” models consistently produce high perplexity, we normalize each model’s perplexity into its z-score within the model. We then compute the mean of z-scored perplexities of the two benchmark signatures within each model and the Spearman correlation between these two mean lists to represent the signature-level overlap, aligning with performance level results and indicating models’ relative relation of perplexity and skill familiarity on the signature. Refer to Appendix A.5.1 for a formalized walk-through.
4 Results
Our experiments are conducted on 32 models and 88 benchmarks, including many of the most widely used ones. We extract benchmark signatures from the open dataset RedPajama (Weber et al., 2024). See Appendix A.5 for the full details of the experimental setup.
4.1 Signatures can better distinguish benchmarks than semantics and model performance
We first examine how the overlap distribution looks across three levels, as illustrated in Figure 4. To minimize inductive bias, we assign broader categories to these benchmarks using the official labels from MMLU (Hendrycks et al., 2021), Big-Bench Hard (Suzgun et al., 2022), ifeval benchmark (Zhou et al., 2023), and MBPP (Austin et al., 2021). In signature overlap (panel a), on the left, we compare within-category overlap against the average cross-category overlap. To reduce the impact of benchmark category size, we ensure each category pair is weighted equally. We then use the mean of cross-category overlaps to represent the overall cross-category overlap and apply this consistently throughout the paper. We observe that overlap is higher within certain categories such as reasoning, science, and social science knowledge, which is expected: benchmarks designed around the same high-level intent tend to align among subtasks, whereas pairs such as chemistry vs. history benchmarks overlap far less. Within the humanities, language and world models, overlaps are generally lower than in cross-category comparisons. A closer look at these benchmarks suggests that these lower similarities arise because they focus on diverse cultural contexts (for example, world models that assess understanding of culture-specific phenomena) and rely on processing across multiple linguistic dimensions. Furthermore, within a category, certain benchmarks under different subtasks align more strongly than others. This forms a dense “red clique,” identified by extracting the maximum clique from the overlap graph. We highlight these highly aligned subtasks on the right side. For panel (b) semantic overlap and panel (c) performance overlap, in contrast, these analyses show much weaker discriminative ability. Semantic overlap scores remain in a narrow range (typically 0.1–0.4) regardless of whether benchmarks come from the same or different categories. Conversely, performance-level overlap is almost universally high, suggesting that model performance and the semantic meaning of questions are less sensitive to category boundaries and obscure finer-grained, underlying associations between benchmarks.
(a) Signature-level analysis
(b) Performance-level analysis
(c) Semantic-level analysis
Figure 4: Three levels of benchmark relation analysis. The signature-level analysis demonstrates substantially stronger discriminative ability compared to both semantic- and performance-level analyses. All heatmaps are presented using a consistent color range from -1 to 1, and panels b and c share the same row and column indices articulated in panel a.
At the semantic level, text embedding models such as MPNet capture surface-level similarity in how humans perceive benchmark questions (Morris et al., 2023). These representations are highly dependent on the specific descriptive intention behind a question, however, meaning the overlap remains superficial and does not reflect the underlying abilities being evaluated. In other words, identical questions do indicate overlapping benchmarks, but different questions do not necessarily indicate non-overlapping ones in terms of underlying ability. At the performance level, while some overlap was initially observed, it quickly became clear that this too fails to meaningfully separate categories. In fact, performance-level results show strong segregation: model behaviors on certain cross-category benchmarks are as closely aligned as they are within categories (evident in several segregated red areas not on the diagonal). When we examine these unexpectedly high alignments, we find that they occur within the same broad benchmark families (e.g., MMLU or BigBenchHard) or under the same question format (e.g., True/False versus multiple-choice questions). This benchmark-orthogonal effect is even stronger than within-category overlaps - that is, MMLU history aligns more closely with MMLU chemistry than with another history benchmark. This underscores the limitations of relying on performance alone and highlights deep issues in current benchmark agreement tests. Several factors could explain this pattern: the bias may stem from post-training fine-tuning, and it could also reflect contamination of the training data, where exposure to one evaluation within a benchmark family increases the likelihood of exposure to others, thereby inflating performance correlations. Another explanation lies in model capabilities: when a model is tested for a single ability, the evaluation inevitably involves a combination of multiple common skills - at a minimum reading, instruction following, and comprehension, among others. This overlap makes behavioral alignment a less distinguishable measure.
4.2 The evaluation bias from model family and question types is resolved by the signature
Grouping the result in Figure 4 panel b, we observe that the red areas are concentrated within the same benchmark family and question format as shown in Figure 1 right panel, where two red areas are exactly two benchmark families or question formats. We calculated pairwise correlations between benchmarks both within and across families and question formats. Since each family or format contains a highly diverse set of benchmarks - essentially covering everything - we would expect within-family/format overlaps to be quite low, showing little difference from cross-family/format overlaps. Consistent with this expectation, the signature-level analysis reveals statistically insignificant tiny differences based on the Mann–Whitney U test, yielding results around 0. This aligns with intuition, as the signature provides a good approximation of the true overlap and variation. In contrast, the performance-level analysis shows a large value of overlap (around 0.8) and a statistically significant increase in within-family/format overlap. Our results show that LLM performance may be more related to surface-level aspects of benchmarks such as question format, suggesting both that generalization and knowledge-propagation in LLMs are limited and that current evaluation may be underestimating peak performance because of conflation of performance and competence.
Figure 5: Biases (within/between families; same/diff. formats) are well addressed by the signature.
4.3 Signatures inform benchmark design and LLM capacity space
As shown in Figure 1, we compare overlaps across design functions. Several patterns emerge. First, we observe significant overlaps that align with intuition. For example, math and logic correlate at 0.21, which is close to the average within-function overlap of 0.285 and far above the average cross-function overlap of 0.087. This makes sense: solving a math problem often requires logical reasoning, and vice versa. More broadly, logic, instruction following, language, math, and world modeling (largely cultural benchmarks) form a cluster of interconnected abilities. Coding appears far less entangled with other functions. Its low cross-function overlap suggests that coding benchmarks are comparatively “clean,” in the sense that success relies more specifically on coding competence and less on auxiliary abilities. This distinctiveness might arise because coding requires highly specialized pretraining corpora such as GitHub.
There are two broad perspectives for interpreting these results. If we optimistically assume that benchmarks faithfully measure what they claim, then the observed overlaps reveal a genuine interdependence of cognitive abilities. In this view, benchmarks are not “leaky,” but rather reflect the multifaceted nature of capacity like math and logic. From this perspective, overlap is not noise, but evidence of underlying LLM and human capacity entanglement - the interconnected capacity space - an often-discussed but previously difficult-to-formalize concept. Alternatively, the overlaps may expose a misalignment between what benchmarks intend to measure and what they actually capture. This interpretation suggests that benchmarks are “leaky” in undesirable ways, inadvertently testing skills outside their stated domain. For example, even if math and logic are highly related, their overlap should theoretically remain lower than within-math or within-logic overlap. Yet, Figure 1 shows cases where cross-function overlap exceeds within-function overlap - for instance, between instruction following and logic. This could imply that either within-function overlap is underestimated (due to poorly aligned benchmark design and execution (Liao et al., 2021)) or that cross-function contamination is stronger than anticipated, undermining the clarity of what each benchmark is supposed to isolate.
5 Final Remarks
Our study introduces benchmark signatures as a principled way to quantify overlaps among LLM benchmarks beyond surface semantics and correlated performance. By grounding benchmark overlap in perplexity patterns from in-the-wild corpora, signatures prove robust against benchmark-orthogonal biases such as model family or question format, while revealing both expected and surprising cross-domain entanglements. These findings point toward a richer and more mechanistic understanding of the LLM capacity space, benchmark validity, and LLM sensitivities. Looking forward, signature-based analyses could extend to finer-grained probes such as LLM layer-level activations when processing benchmarks, or even a form of “benchmark algebra” to systematically identify limitations in the current evaluation landscape.
Appendix A Appendix
A.1 Related Literature
Benchmark Categorization and Overlap: Benchmarks are central to model evaluation. Two simple metrics capture their utility: signal, a benchmark’s ability to reliably distinguish better models from worse ones, and noise, a benchmark’s sensitivity to randomness (Heineman et al., 2025). Recently, researchers have begun to ask how comparable benchmarks with similar intent actually are. This is commonly studied through Benchmark Agreement Testing (BAT), where new benchmarks are validated against established ones using agreement metrics (e.g., rank correlation) (Perlitz et al., 2024). Such analyses have led to concerns that the community may be producing too many benchmarks. For example, Liu et al. (Liu et al., 2021) examined agreement across multiple QA benchmarks and concluded that because agreement was high, additional QA benchmarks were unnecessary. Beyond statistical agreement, some recent works have attempted to qualitatively interpret and categorize benchmarks - for example, as testing logical reasoning or commonsense reasoning - though often without running agreement tests either within or across these categories (Ni et al., 2025). Another emerging line of inquiry asks what capabilities are still missing from current benchmark suites. Miller and Tang (Miller & Tang, 2025), for instance, examine how people commonly use LLMs for summarization, technical assistance, reviewing work, data structuring, generation, and information retrieval, and assess the extent to which existing benchmarks cover these capabilities. Their findings reveal significant gaps in coverage of benchmarks across categories.
Signal Extraction from Pretraining Data: A growing body of work investigates how information extracted from pretraining corpora can inform data selection and model evaluation, even building benchmarks automatically. A central insight is that LLM losses on pretraining texts are often correlated with downstream benchmark performance, suggesting that simple loss-performance correlation coefficients can be effective signals for identifying high-quality pretraining data from in-the-wild corpus (Thrush et al., 2025; Hoffmann et al., 2022). Validation loss is thus frequently used as a proxy for model generalization (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022), and with more recent evidence showing that such correlations persist across architectures and training settings (Poli et al., 2023). One line of research focuses on efficient, low-cost methods for understanding and filtering signals, for instance lightweight approaches using surface-level heuristics (n-gram overlap (Xie et al., 2023) or semantic-level similarities (Everaert & Potts, 2023)), enabling scalable filtering of massive corpora. Thrush et al., (Thrush et al., 2025) proposed an orthogonal approach for data selection centered around estimates of perplexity-benchmark correlations. We build on these ideas to construct benchmark signatures by mining predictive tokens of LLM performance from large-scale pretraining corpora, in order to address challenges in meta-evaluation - the evaluation of LLM evaluations, e.g., how overlapping they are.
A.2 Comparisons between token-, chunk-, and document-level perplexity
From fine to coarse granularity, we consider token-, chunk-, and document-level perplexities. At the document level, we evaluate the model on an entire pretraining document and take the mean across all text chunks that fit within the model’s context window (part of the document). At the chunk level, we split documents into fixed-length windows (30 pieces, using spaces as separators) and compute perplexity as the average over all tokens within each window. At the token level—the finest granularity with the least inductive bias—we use token-wise perplexities from documents to capture the model’s intrinsic uncertainty. Concretely, we form a window by taking the target token with its up-to-30 preceding pieces (using spaces as separators) as context, then record only the last token’s perplexity as the feature. This ensures the token is conditioned on its preceding context rather than treated in isolation. As shown in Table 1, the token level exhibits the greatest standard deviation and interquartile range, as well as more pronounced extreme gaps in majority cases compared to the chunk and doc levels. This wider dispersion indicates that extreme values are more visible and significant at the token level, making it a natural choice for feature selection. Token-level signatures balance the strongest predictive power (both positive and negative relations) of highly informative tokens, and they exhibit high variance of predictive power, as captured by the deviations. By focusing on the token level, we are able to highlight more prominent signals, whereas aggregation at the chunk or document level tends to smooth out these extremes.
Benchmark Level Std IQR Max–Q99 Q01–Min
Gsm8k Chunk 30.30 44.63 43.37 36.63
Gsm8k Doc 19.39 26.50 23.25 18.75
Gsm8k Token 36.53 54.11 40.63 27.37
Mbpp Chunk 29.83 43.05 87.62 27.05
Mbpp Doc 14.49 18.78 32.44 8.67
Mbpp Token 36.76 50.60 39.60 39.20
Mmlu Chunk 29.31 41.18 64.00 30.36
Mmlu Doc 18.11 27.68 35.16 9.68
Mmlu Token 38.05 51.90 52.00 42.29
Truthfulqa Chunk 30.58 41.80 48.00 29.20
Truthfulqa Doc 17.67 22.75 25.00 13.00
Truthfulqa Token 36.57 53.00 49.20 35.40
Table 1: Summary of Thrush coefficient distributions across four benchmarks. Columns report standard deviation (Std), interquartile range (IQR), and tail gaps of Max-Q99 and Q01-Min are defined as the distance from the maximum to the 99th percentile (Max–99th) and from the 1st percentile to the minimum (1st–Min). Across 16 targets (4 measures and 4 benchmarks) token-level values achieved 12 wins.
A.3 Conditions for Sure Independence Screening (SIS)
Sure Independence Screening (SIS) is a powerful statistical tool for feature selection in ultra-high dimensional settings, offering a “sure screening property” that guarantees the retention of truly informative features with high probability under specific conditions (Fan & Lv, 2008). In this section, we elaborate on how the key theoretical assumptions underlying SIS are plausibly met within our problem context of mining benchmark signatures from token-level perplexities.
  1. 1.
    Ultra-High Dimensionality: Our problem inherently operates in an ultra-high dimensional regime, where the number of ”in-the-wild” tokens ($d$, scaling to billions) vastly exceeds the number of language models ($m$, typically in the tens). Specifically, we have $\log(d)>m$, which far exceeds the standard $d>m$ high-dimensional definition. This extreme disparity makes full multivariate regression computationally intractable, underscoring the necessity of an efficient screening step like the one we employ.
  2. 2.
    Sparsity: The “Sparsity” assumption (our [A1]) posits that only a small fraction of the $d$ tokens are truly informative for predicting LLM benchmark performance. Our empirical observations of the correlation coefficient distributions (e.g., Figure 3) directly support this. The distributions show a strong concentration around zero, indicating that most tokens have little to no marginal predictive power. The presence of thin but distinct tails also suggests that a small subset of tokens exhibits strong correlations, aligning with the idea that specific linguistic phenomena (represented by these tokens) drive performance on a given benchmark.
  3. 3.
    Minimum Signal Strength: SIS requires that the true predictive signals (i.e., the tokens with non-zero effects on benchmark performance) are not arbitrarily weak. In our context, this translates to these important tokens having sufficiently strong marginal correlations to stand out from the noise. Our use of token-level perplexities, which directly reflect an LLM’s familarity of specific linguistic patterns, suggests that truly important tokens would indeed manifest as strong signals. The robust, rank-based correlation coefficients we employ (Thrush and Pre-select) are also well-suited to detect such signals, as they are less sensitive to outliers and distributional peculiarities that might obscure signals when using less robust measures.
  4. 4.
    Limited Pathological Multicollinearity: A critical condition for basic SIS is that the multicollinearity between important features and unimportant ones should not be so severe that it masks the marginal signal of truly predictive tokens (e.g., the suppressor variable scenario). While token perplexities can exhibit correlations (e.g., highly similar tokens or tokens from common linguistic constructs), it is less probable that a truly causal token’s signal would be perfectly canceled out by others at a marginal level. Benchmarks typically probe specific abilities, which are likely associated with a distinct, though perhaps overlapping, set of “signature” tokens. The vast and diverse nature of “in-the-wild” tokens also means that while many tokens might be highly correlated, there are many more effectively independent ones. More importantly, the core objective of our work is to identify benchmark signatures as a specific and parsimonious set of tokens. If a token’s marginal signal is entirely masked, it might suggest its contribution is highly redundant with other tokens that do have a strong marginal signal, or that its unique contribution is extremely weak – in which case, its exclusion from the initial screening might not significantly harm the final signature’s predictive power.
A.4 Technical Details
A.4.1 Semantic-level Bootstrapped Similarity Calculation
Algorithm 2 Get Pairwise Similarity Matrix
0: A list of benchmarks $\mathcal{B}=\{B_{1},\ldots,B_{n}\}$; Embedding model $E$ (e.g. a sentence transformer); Number of bootstrap replicates $k$.
0: An $n\times n$ similarity matrix $S$.
[…]
3:for $i=1$ to $n$ do
4:  for $j=i$ to $n$ do
5:   if $i=j$ then
6:    $S_{i,j}\leftarrow 1.0$
7:   else
8:    $s\leftarrow\text{getSimScore}(B_{i},B_{j},E,k)\triangleright$ Call Algorithm 3
9:    $S_{i,j}\leftarrow s$
10:    $S_{j,i}\leftarrow s\triangleright$ Matrix is symmetric
11:   end if
12:  end for
13:end for
14:return $S$
Algorithm 3 Bootstrapped Similarity Score Calculation (getSimScore)
0: Benchmarks AA and $B$; Embedding model $E$; Bootstrap replicates $k$.
0: A single similarity score $\text{sim}_{A,B}$.
[…]
2:if $|A|<|B|$ then
3:  $S\leftarrow A$; $L\leftarrow B$ $\triangleright$ $S$ is the smaller, $L$ is the larger
4:else
5:  $S\leftarrow B$; $L\leftarrow A$
6:end if
[…]
15:for $b=1$ to $k$ do
16:  $L^{\prime}\leftarrow\text{sample}(L,n_{s},\text{replace}=\text{False})$
17:  $\text{text}_{L}\leftarrow\text{concatenate all questions in }L^{\prime}$
18:  $\text{text}_{L}^{\prime}\leftarrow\text{truncate}(\text{text}_{L},\ell)$
19:  Append $\text{text}_{L}^{\prime}$ to $\mathcal{T}_{L}$
20:end for
[…]
25:return $\text{sim}_{A,B}$
A.4.2 AIC Stepwise Forward Selection Algorithm
Algorithm 4 Selecting Salient Tokens for Benchmark $j$
0:  Perplexity feature matrix $\mathbf{P}$; performance vector $\mathbf{y}_{:,j}$; tail fraction $\alpha=0.01$; tolerance $\delta\geq 0$
0:  Salient Token set $\mathcal{T^{\prime}}_{j}$
1:  Preselection via Thrush Correlation
2:  for $\ell=1$ to $d$ do
3:   $\rho_{\ell}\leftarrow\mathrm{ThrushCorr}(\mathbf{P}_{:,\ell},\mathbf{y}_{:,j})$
4:  end for
5:  $T^{+}\leftarrow$ indices of the top $\alpha d$ values of $\rho_{\ell}$ $\triangleright$ most positively correlated
6:  $T^{-}\leftarrow$ indices of the bottom $\alpha d$ values of $\rho_{\ell}$ $\triangleright$ most negatively correlated
7:  $T\leftarrow\mathrm{Shuffle}(T^{+}\cup T^{-})$ $\triangleright$ candidate feature set
8:  Forward Selection with AIC (on $T$)
9:  $S\leftarrow\emptyset$;   $A^{\star}\leftarrow+\infty$
10:  while $T\setminus S\neq\emptyset$ do
11:   for each $\ell\in(T\setminus S)$ do
12:    $A(\ell)\leftarrow\mathrm{AIC}\!\left(\mathrm{Fit}(\mathbf{y}_{:,j}\sim\mathbf{P}_{:,\;S\cup\{\ell\}})\right)$
13:   end for
14:   $\ell^{\star}\leftarrow\arg\min_{\ell\in(T\setminus S)}A(\ell)$;   $A_{\text{new}}\leftarrow A(\ell^{\star})$
15:   if $A_{\text{new}}<A^{\star}-\delta$ then
16:    $S\leftarrow S\cup\{\ell^{\star}\}$;   $A^{\star}\leftarrow A_{\text{new}}$
17:   else
18:    break $\triangleright$ no further AIC improvement
19:   end if
20:  end while
21:  $\mathcal{T^{\prime}}_{j}\leftarrow S$
22:  return $\mathcal{T^{\prime}}_{j}$
A.5 Experiment Setup
Overview: Our chosen benchmarks span diverse domains such as knowledge (business, humanities, social sciences, science and engineering, medicine), mathematics, coding, reasoning, language, culture and world knowledge, logic, and instruction following. We choose 32 widely-used language models (see the list below). We extract benchmark signatures from the open dataset RedPajama (Weber et al., 2024), which contains large-scale textual data across multiple domains, including CommonCrawl, C4, GitHub, arXiv, Books, Wikipedia, and StackExchange, used for pretraining LLMs, making it a strong source of in-the-wild data for mining benchmark signatures. We take the standard approach, using vLLM (Kwon et al., 2023) for facilitating perplexity extraction and llm-evalution-harness (Gao et al., 2024) for evaluation across benchmarks and models such that all evaluations are under the same condition.
A.5.1 Experiment Walkthrough
As discussed, we measure perplexity at three granularities - token, chunk, and document levels (from fine to coarse). The segmentation procedure for each is detailed in §A.2. We ultimately focus on the token level because it provides the clearest view of prominent signals for the pre-filtering stage.
Preprocessing RedPajama
We use the 1B-token RedPajama variant to balance scale and computational cost. For token-level segmentation, we split the corpus on whitespace into pieces. For each piece, we prefix up to the preceding 30 pieces as left context and record the last token’s perplexity conditioned on that context. This yields an initial pool on the scale of billions of token-level contexts ($d\approx 8.45\times 10^{9}$). To reduce noise or in-the-wild text, we uniformly downsample by a factor of $1/50$, yielding approximately $1.69\times 10^{7}$ instances.
Feature Matrix Construction
Using the vLLM setup described in §A.5.4, we evaluate 32 models on the token contexts and extract token-level perplexities, forming the covariate (feature) matrix
$\mathbf{P}\in\mathbb{R}^{32\times 1.69\times 10^{7}},$
with rows indexed by models and columns by token instances.
Performance Matrix Construction
In parallel, we compute model performance on a series of benchmarks and subfields using the lm-evaluation-harness (details in §A.5.5). Let
$\mathbf{Y}\in\mathbb{R}^{32\times 88}$
denote the performance matrix (models $\times$ benchmarks/subfields). For each benchmark $B_{j}$, the vector $\mathbf{y}_{:,j}$ is the performance vector for $B_{j}$ across all 32 models.
Filtering with Thrush
For each benchmark $B_{j}$, we compute the Thrush rank correlation between the entire feature matrix $\mathbf{P}$ and the performance vector $\mathbf{y}_{:,j}$. This produces a distribution of Thrush scores over token features. We retain the top 1% and bottom 1% features (by score) and concatenate these extremes into a benchmark-specific subset of columns from $\mathbf{P}$ for downstream modeling.
AIC Step-Forward Feature Selection
Finally, for each benchmark $B_{j}$, we fit a multivariate linear model on the preselected features using step-forward selection with the Akaike Information Criterion (AIC) as the objective. Starting from an empty model, we iteratively add the feature that most improves AIC and stop when no further improvement is possible (tolerance =0=0). The resulting selected set constitutes the most predictive in-the-wild token features for $B_{j}$. Across benchmarks, the selected set size varies but typically has $\sim 30$ features.
Signature and Comparison
Consider two benchmark signature vectors, $S_{1}$ and $S_{2}$, each consisting of several context pieces (30 pieces separated by spaces) plus the salient token. For each benchmark we acquire around 30 such non-overlapping salient tokens. We evaluate these signatures with 32 models, which read their respective pre-contexts and compute last-token perplexities. If the models exhibit similar levels of perplexity for both signatures, this strongly suggests that the two benchmarks align. We normalize each model’s perplexity values into their $z$-score within the model. For each model, we then compute the mean of the $z$-scored perplexities for the two benchmark signatures. Finally, we calculate the Spearman correlation ($\rho_{s}$) between these two mean vectors to represent signature-level overlap.
A.5.2 Datasets
Benchmark family Benchmarks
mmlu business_ethics; marketing; management; professional_accounting; high_school_european_history; jurisprudence; humanities; prehistory; professional_law; world_religions; high_school_us_history; formal_logic; high_school_world_history; international_law; logical_fallacies; moral_disputes; moral_scenarios; philosophy; anatomy; clinical_knowledge; college_medicine; human_aging; medical_genetics; nutrition; professional_medicine; virology; miscellaneous; other; abstract_algebra; astronomy; college_biology; college_chemistry; college_physics; conceptual_physics; elementary_mathematics; high_school_biology; high_school_computer_science; college_computer_science; college_mathematics; computer_security; electrical_engineering; high_school_chemistry; high_school_mathematics; high_school_physics; high_school_statistics; machine_learning; stem; high_school_geography; high_school_government_and_politics; high_school_macroeconomics; high_school_microeconomics; high_school_psychology; public_relations; us_foreign_policy; econometrics; human_sexuality; professional_psychology; security_studies; social_sciences; sociology
bbh global_facts; hyperbaton; snarks; web_of_lies; word_sorting; disambiguation_qa; salient_translation_error_detection; boolean_expressions; dyck_languages; geometric_shapes; multistep_arithmetic_two; navigate; object_counting; reasoning_about_colored_objects; formal_fallacies; logical_deduction_five_objects; logical_deduction_seven_objects; logical_deduction_three_objects; penguins_in_a_table; temporal_sequences; tracking_shuffled_objects_five_objects; tracking_shuffled_objects_seven_objects; tracking_shuffled_objects_three_objects; causal_judgement; movie_recommendation; ruin_names; date_understanding; sports_understanding
Table 2: Benchmarks in MMLU and BBH.
Format Benchmarks
multi-choice questions business_ethics; marketing; management; professional_accounting; high_school_european_history; jurisprudence; humanities; prehistory; professional_law; world_religions; high_school_us_history; formal_logic; high_school_world_history; international_law; logical_fallacies; moral_disputes; moral_scenarios; philosophy; anatomy; clinical_knowledge; college_medicine; human_aging; medical_genetics; nutrition; professional_medicine; virology; miscellaneous; other; abstract_algebra; astronomy; college_biology; college_chemistry; college_physics; conceptual_physics; elementary_mathematics; high_school_biology; high_school_computer_science; college_computer_science; college_mathematics; computer_security; electrical_engineering; high_school_chemistry; high_school_mathematics; high_school_physics; high_school_statistics; machine_learning; stem; high_school_geography; high_school_government_and_politics; high_school_macroeconomics; high_school_microeconomics; high_school_psychology; public_relations; us_foreign_policy; econometrics; human_sexuality; professional_psychology; security_studies; social_sciences; sociology
true or false boolean_expressions; causal_judgement; formal_fallacies; navigate; sports_understanding; web_of_lies
Table 3: Classification of benchmarks by question format.
A.5.3 Models Used
The models we used in this study are summarized in Table 4. Each entry cites the official paper if available, otherwise the model card.
Model (HF Repository) Citation
meta-llama/Llama-3.1-8B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-1B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-3B-Instruct Grattafiori et al. (2024)
google/gemma-3-4b-it Team (2025a)
google/gemma-3-12b-it Team (2025a)
google/gemma-3-27b-it Team (2025a)
mistralai/Mistral-7B-Instruct-v0.3 Jiang et al. (2023)
deepseek-ai/deepseek-llm-7b-chat DeepSeek-AI et al. (2024)
Qwen/Qwen3-0.6B Team (2025b)
Qwen/Qwen3-1.7B Team (2025b)
Qwen/Qwen3-4B Team (2025b)
Qwen/Qwen3-8B Team (2025b)
tiiuae/falcon-rw-1b Penedo et al. (2023)
EleutherAI/pythia-1b Biderman et al. (2023)
EleutherAI/pythia-6.9b-v0 Biderman et al. (2023)
EleutherAI/pythia-12b-deduped Biderman et al. (2023)
mosaicml/mpt-7b-instruct Team (2023)
microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024a)
microsoft/Phi-4-mini-instruct Microsoft et al. (2025)
microsoft/Phi-4 Abdin et al. (2024b)
01-ai/Yi-1.5-9B-Chat AI et al. (2025)
01-ai/Yi-1.5-6B-Chat AI et al. (2025)
mistralai/Ministral-8B-Instruct-2410 Jiang et al. (2024)
openai-community/gpt2-medium Radford et al. (2019)
openai-community/gpt2-large Radford et al. (2019)
openai-community/gpt2-xl Radford et al. (2019)
zai-org/chatglm3-6b GLM et al. (2024)
zai-org/glm-4-9b-chat-hf GLM et al. (2024)
zai-org/codegeex4-all-9b Zheng et al. (2023)
allenai/OLMo-2-1124-13B-Instruct OLMo et al. (2024)
allenai/OLMo-2-1124-7B-Instruct OLMo et al. (2024)
allenai/OLMo-2-0425-1B-Instruct OLMo et al. (2024)
Table 4: Models used in this study. Multiple variants may share the same citation.
A.5.4 Inference with vLLM
We use vLLM (Kwon et al., 2023) for facilitating perplexity extraction. Specifically, we run all inference in offline mode with tensor parallelism across 2 GPUs to maximize throughput and data parallelism across 1 GPU. The model weights are cached locally to reduce repeated I/O overhead, and inference is performed in batches of prompts to further amortize the computation cost. For each sequence, we extract per-token log probabilities, from which we compute negative log-likelihood and perplexity metrics. All outputs are aggregated into parquet for downstream analysis. The GPUs we used in this computation are $2\>\times$ Nvidia A100 (80GB).
A.5.5 Benchmark Evaluation
For benchmark evaluation we use llm-evalution-harness (Gao et al., 2024). Each benchmark involves different task formats, and we adopt the standard metrics defined in the harness to ensure comparability. For MMLU, which consists of multi-choice questions across 57 academic subjects, we report accuracy, i.e., the proportion of questions with correctly selected options. For MBPP, a code generation benchmark, we evaluate using pass@1, the fraction of problems solved correctly on the first attempt, based on unit test execution. For BBH, which is a collection of heterogeneous tasks (multiple-choice, binary classification, and completion), we follow the harness in applying the canonical metric for each subtask. We use accuracy for multiple-choice and true/false items, and exact match for sentence completions. For IFEval, which tests instruction-following, we adopt the harness’s compliance accuracy, quantifying the percentage of model responses that satisfy the explicit constraints in the prompt. These heterogeneous metrics reflect the intended difficulty and modality of each benchmark, and together provide a broad view of model capability.